### v1

# What’s under the hood: Investigating Automatic Metrics on Meeting Summarization

Abstract
Meeting summarization has become a critical task considering the increase in online interactions. While new techniques are introduced regularly, their evaluation uses metrics not designed to capture meeting-specific errors, undermining effective evaluation. This paper investigates what the frequently used automatic metrics capture and which errors they mask by correlating automatic metric scores with human evaluations across a broad error taxonomy. We commence with a comprehensive literature review on English meeting summarization to define key challenges like speaker dynamics and contextual turn-taking and error types such as missing information and linguistic inaccuracy, concepts previously loosely defined in the field. We examine the relationship between characteristic challenges and errors by using annotated transcripts and summaries from Transformer-based sequence-to-sequence and autoregressive models from the general summary QMSum dataset. Through experimental validation, we find that different model architectures respond variably to challenges in meeting transcripts, resulting in different pronounced links between challenges and errors. Current default-used metrics struggle to capture observable errors, showing weak to mid-correlations, while a third of the correlations show trends of error masking. Only a subset reacts accurately to specific errors, while most correlations show either unresponsiveness or failure to reflect the error’s impact on summary quality.
What’s under the hood: Investigating Automatic Metrics on Meeting Summarization
Frederic Kirstein1,2, Jan Philip Wahle1, Terry Ruas1, Bela Gipp1 1Georg-August-Universität Göttingen, Germany 2kirstein@gipplab.org
1 Introduction
With the surge in remote collaboration, there’s an amplified need for proficient meeting summarization Mroz et al. (2018); Pratama et al. (2020), beneficial for participants and non-attendees. Despite advancements in addressing current meeting challenges like contextual turn-taking and discourse structure, existing evaluation metrics often inadequately reflect these Rennard et al. (2023); Kumar and Kabiri (2022). The widespread default use of the ROUGE metric Lin (2004) has shown limitations Liu and Liu (2008); Cohan and Goharian (2016); Fabbri et al. (2021). While the research community acknowledges the potential of newer metrics, such as BERTScore Zhang et al. (2020b), their efficacy in meeting summarization remains uncharted. This is a crucial shortcoming, as the high abstraction level, low extraction rate, and the requirement for complex reasoning power of meeting summarization present new challenges to automatic evaluation metrics, which need to be explored Gao and Wan (2022). Our study aims to investigate how commonly used automatic metrics correlate with human annotations, exploring the inherent qualities these metrics encapsulate in their scores. To unify the fragmented understanding of challenges posed by meeting summarization and the typical errors that manifest when these challenges are unmet. Using the QMSum dataset Zhong et al. (2021), we employ experts to annotate challenges in meeting transcripts and errors in summaries. Various models, including domain-standard encoder-decoder architectures and notable autoregressive models, generated these summaries. Our methodological approach enables us to draw correlations between the challenges and errors, the annotated errors, and a set of eight automatic metrics. The insights derived reveal shortcomings in the current automatic metrics employed for meeting summarization evaluation: Structural disorganization errors are often penalized, reflecting human assessments, whereas hallucination errors tend to be rewarded. While adept at penalizing omissions, the ROUGE Lin (2004) metric struggles to discern the nuances in error impact on quality. Strikingly, around a third of metrics-error combinations either overlook or reward errors. For instance, Perplexity leans towards incorrect references, and Lens tends to favour structural disorganization. This underscores a pressing demand for refined evaluation methods in meeting summarization. Our contributions are threefold:
  • We conduct a comprehensive literature review to identify and unify the challenges specific to English meeting summarization and the types of errors that commonly occur in model-generated summaries.
  • We build the first direct correlations between the intrinsic challenges and observable errors they induce for encoder-decoder and autoregressive model architectures, providing a framework for tracking their impact.
  • We rigorously evaluate the efficacy of nine prevalent automatic metrics, uncovering what they accurately measure, neglect, and their sensitivity to varying error severities, all corroborated by human annotations.
We will release the codebases, annotations and guidelines to aid continued research in this domain.
2 Methodology
We conducted a focused literature review to identify challenges and observable errors when failing these challenges in English meeting summarization, specifically targeting papers published after the seminal 2019 introduction of the transformer-based BART Lewis et al. (2020) model, which left a lasting mark in the field. This further allows us to target the challenges in meeting summarization with our survey as they emerge and pertain to the currently used Transformer-based models (e.g., coreference resolution and spoken language). Besides this temporal filter, we set inclusion criteria to prioritize relevance over historical comprehensiveness and consider works that either introduce novel methodologies or contribute to defining challenges, with a preference for those employing transformer architectures. Exclusion criteria eliminate tangential studies, highly specialized papers not addressing broader challenges, multi-lingual and multi-modal approaches, and duplicates. We considered publications from 2023 irrespective of their citation index, acknowledging their emergent influence. The survey thereby orients on the PRISMA checklist to prevent bias.
Our primary database was Google Scholar , chosen for its comprehensive coverage of both extractive and abstractive methodologies and advanced query capabilities. While ACL Anthology , Web of Science , and DBLP were assessed, Google Scholar yielded a more nuanced spectrum of relevant literature, conforming to our review scope spanning from 2019 to September 2023.
A two-stage search strategy was employed: The initial phase involved broad queries combining "meeting summarization" with adjunctive terms such as challenge, literature review, or survey, combined through a logical operation. Subsequent refinement employed challenge-related keywords and synonyms retrieved from the review literature, ensuring alignment with our research focus. Papers were subsequently ranked based on their Google Scholar metrics, with the top 100 from each query considered. The corpus thus aggregated served as the primary source for our analysis.
From the initial pool of 300 papers, we rigorously applied our criteria to select 18 core articles addressing challenges directly and an additional 40 that discuss challenges and observable errors indirectly through methodological contributions. Each shortlisted paper underwent detailed data extraction to identify relevant challenges, methodologies, foundational models, and metrics, which were then collated in a structured format for cross-validation. This allowed us to construct a comprehensive view of current challenges and observable errors in English meeting summarization.
3 Literature Review
3.1 Challenges
This section summarizes key challenges in meeting summarization identified in the literature and their implications on the summarization process.
Spoken language. Handle colloquialisms, domain-specific terminology, and various forms of linguistic noise, such as false starts, repetitions, and filler words Koay et al. (2020); Kumar and Kabiri (2022); Feng et al. (2022); Antony et al. (2023); Zhang et al. (2021), which can appear due to the spoken language nature of meetings. This can affect the accuracy and clarity of the generated summaries.
Speaker dynamics. Accurately distinguish and track different speakers, their utterances, and specific roles (e.g., project manager, applicant), particularly when roles are topic-dependent Gu et al. (2022); Rennard et al. (2023); Shinde et al. (2022); Feng et al. (2022); Antony et al. (2023); Kumar and Kabiri (2022); Khalifa et al. (2021). Failing to do so can introduce biases and result in incomplete or flawed summaries leaving out noteworthy elements.
Coreference. Manage the resolution of references to other speakers and previous actions to ensure coherent and complete summarization Liu et al. (2021). Inaccurate handling can lead to ambiguous or incomplete summaries lacking contextual richness.
Discourse structure. Understand and track the inherent high-level structure and flow of a meeting Li et al. (2023) throughout the different meeting phases. These phases may refer to multiple topics Feng et al. (2022), e.g., during an argumentation, or swiftly shift from one topic to another Shinde et al. (2022); Feng et al. (2022). Failure to consider this hierarchical structure can result in summaries that are either incomplete or lack coherence.
Contextual turn-taking. Capture the evolving local dynamics of a meeting as it progresses through multiple speaker turns. The task involves accurately processing shifts in discourse, complicated by interruptions, repetitions, and redundancies Ma et al. (2022); Gu et al. (2022); Zhang and Zhao (2021); Zhang et al. (2021); Shinde et al. (2022). Inadequate capture can result in misleading or shallow summaries.
Implicit context. Account for unspoken or implied context, such as tacit organizational knowledge or prior discussions that are not explicitly referred to. Neglecting this can produce misleading or shallow summaries Xu et al. (2022).
Low information density. Identify key points when salient information is sparse and unevenly distributed, especially relevant in decision-making scenarios Rennard et al. (2023); Zhang et al. (2021). This can affect the depth and level of salience in the resulting summary.
Data scarcity. The lack of diverse, high-quality real-world scenario training samples hampers model training Rennard et al. (2023); Kumar and Kabiri (2022); Jacquenet et al. (2019).
Long transcripts. Process long transcripts, which can result from longer meetings, due to the quadratic computational cost of transformer-based models, leading to efficiency issues with increasing dialogue length Rennard et al. (2023); Kumar and Kabiri (2022); Shinde et al. (2022); Feng et al. (2022); Jacquenet et al. (2019); Zhang et al. (2021). Only processing a sub-part of a meeting might lead to the exclusion of salient information and therefore result in an incomplete or wrong summary.
Heterogeneous meeting formats. Different types of meetings necessitate distinct summarization approaches Rennard et al. (2023). Failure to adapt can result in summaries that miss key points or include irrelevant information
3.2 Error Types
Errors in summaries arise when challenges aren’t correctly addressed. We’ve collated error types from the literature into six principal categories, which were observed in the summaries generated during our experiments as detailed in Section 5.
Missing information. This error involves missing information from the meeting, such as significant decisions or actions Zou et al. (2023); Chen and Yang (2020a). We consider total omission, where relevant topics or points discussed are entirely absent from the summary, and insufficient detail, where the summary mentions a salient topic but does not capture its depth or the detailed discussion from the original meeting.
Redundancy. The summary contains repeated or redundant information, impacting brevity and clarity Chen and Yang (2020a). Such repetitions can manifest in different ways: through reiterated key points, overuse of individual words, or the duplication of entire phrases.
Wrong references. The model misattributes statements, opinions, or actions to incorrect meeting participants or omits their mention altogether Chen and Yang (2020a).
Incorrect reasoning. The model draws conclusions that are not supported by the discussions in the meeting Chen and Yang (2020a).
Hallucination. The model produces inconsistencies, such as incorrect dates, names, or locations, not aligned with the meeting content Wang et al. (2022a); Ji et al. (2023). This encompasses intrinsic hallucinations, introducing events not present in the original meeting or contradicting the input, and extrinsic hallucinations, misrepresenting actual events.
Incoherence. The model generates summaries with disjointed logic or flow. This manifests as intra-sentence disconnections and inter-sentence inconsistencies, such as flawed progression and incorrect contrastive expressions Wang et al. (2023).
During our human annotation process, as detailed in Section 4.4, we identified two additional error categories. Notably, our review of the pertinent literature revealed that these categories have not been explicitly defined in existing studies in a way that matches our findings.
Linguistic inaccuracy. The model uses inappropriate, incorrect, or ambiguous language or fails to capture unique linguistic styles. This error type spans issues from adopting unsuitable words from the source to grammar mistakes, as well as employing contextually unclear or ambiguous terms.
Structural disorganization. The model may produce summaries that misrepresent the original order or logic of the meeting’s discourse, misplacing topics or events. This error covers only when the order of stated events is wrong. The inclusion of non-existing events or exclusion of events does not count towards this error.
4 Experimental Framework
Dataset # Meetings # Turns # Speakers avg. Len. of Meet. avg. Len. of Sum.
ICSI 137 535.6 4.0 6007.7 70.5
AMI 59 819 6.3 13317.3 53.7
WPCP 36 207.7 34.1 13761.9 80.5
all (QMSum) 232 556.8 9.2 9069.8 69.6
Table 1: Statistics for the QMSum subsets and the entire QMSum meeting summarization dataset.
4.1 Models
Our primary model is the Longformer Encoder Decoder (LED) Beltagy et al. (2020) given its recognized summarization task performance. LED employs a unique local+global sliding window attention in its encoder and full self-attention in its decoder, accommodating lengthy documents efficiently. For a comprehensive assessment of challenges and errors in the context of the common encoder-decoder architecture, we incorporated DialogLED Zhong et al. (2022) and PEGASUS-X Phang et al. (2022). DialogLED extends LED with dialogue-centric pre-training. In contrast, PEGASUS-X enhances the original PEGASUS Zhang et al. (2020a) for extended inputs using staggered block-local attention. These models have been finetuned on QMSum.
We also ventured into large autoregressive language models for meeting summarization, leveraging GPT3.5 turbo via ChatGPT and Zephyr-7B-$\alpha$, a refined version of Mistral-7B-v0.1111https://huggingface.co/mistralai/Mistral-7B-v0.1. Mistral’s adaptations employ sliding window attention and have gained recent attention for outperforming Llama2 Touvron et al. (2023). These models, exemplifying the trends in language models, excel in zero-shot summarization Bhaskar et al. (2023). To tackle context size limitations, we employed a chunk-based approach Bhaskar et al. (2023), using the prompt "Create a TL;DR of the following meeting chunk", inspired by a Microsoft guideline  [^2] https://www.microsoft.com/en-us/microsoft-365-life-hacks/organization/using-chatgpt-creating-meeting-agendas-minutes-notes.
4.2 Datasets
We employ QMSum Zhong et al. (2021), a specialized, widely established, and currently most comprehensive dataset for query-based multi-domain meeting summarization. It includes a diverse range of meeting transcripts from academic meetings (ICSI), product meetings (AMI), and committee meetings (Welsh Parliament, Canadian Parliament: WPCP). Detailed statistics are available in Table 1.
ICSI Janin et al. (2003) offers transcripts from informal research meetings, showcasing linguistic challenges such as missing words, sentence restarts, and fillers. AMI Mccowan et al. (2005) provides structured meetings with predefined participant roles and agendas yet still captures natural dialogue dynamics like interruptions and non-verbal cues. WPCP presents formal, agenda-driven discussions from UK and Canada Parliament committee meetings, allowing for a more structured summarization task.
SL SD Cor DS CTT IC DD LID overall
challenges 0.84 0.77 0.68 0.79 0.73 0.90 0.87 0.78 0.79
frequency 0.84 0.71 0.78 0.70 0.89 0.91 0.90 0.84 0.82
Table 2: Krippendorff’s alpha for inter-annotator agreement on challenges and their frequencies, with challenges (abbreviations) in top row ordered as in Section 3.1.
MI Red WR IR Hal Inc LI SD overall
Encoder-Decorder 0.74 0.875 0.83 0.87 0.79 0.61 0.86 0.71 0.78
impact (LED) 0.82 0.92 0.85 0.73 0.80 0.89 0.87 0.79 0.83
Autoregressive 0.76 0.87 0.85 0.79 0.77 0.90 0.90 0.71 0.82
Table 3: Krippendorff’s alpha for inter-annotator agreement on errors for encoder-decoder (i.e., LED, DialogLED) and autoregressive models (i.e., GPT-3.5, Zephyr-7B-$\alpha$). For the LED model, the agreement on error impact is also reported. Errors (abbreviations) in the top row are ordered as in Section 3.2.
4.3 Metrics
For aligning human annotation with automatic metrics, we employ a suite of metrics spanning count-based, model-based, and QA-based methodologies, chosen based on their prevalent use in meeting and dialogue evaluation Gao and Wan (2022).
4.3.1 Count-based
ROUGE Lin (2004) is the default-used metric suite that assesses the overlap between n-grams in generated summaries that appear in the reference. Researchers mainly consider unigrams, bigrams, and the longest common sequence to gauge the relevance and accuracy of generated content. BLEU Papineni et al. (2002) evaluates how many n-grams from the reference appear in the generated summary. The score is designed to reflect the precision of the generated text, focusing primarily on lexical similarity to the reference. METEOR Banerjee and Lavie (2005) builds on BLEU by accounting for synonyms, word stems, and sentence structure, offering a holistic assessment of lexical, syntactic, and semantic alignment in precision and recall.
4.3.2 Model-based
BERTScore Zhang et al. (2020b) utilizes a pre-trained BERT model to measure the contextual similarity between the generated and reference texts, reflecting both semantic and syntactic similarity. We report the rescaled  [^3] https://github.com/Tiiiger/bert score/blob/master/journal/rescale baseline.md BERTScore-F score. Perplexity (PPL) measures the uncertainty of a language model in predicting each word in a given text, with lower scores denoting better predictions. It evaluates the model’s grasp of language structure and word occurrence patterns, thereby evaluating the quality and fluency of utterances. We use GPT-2 Radford et al. (2019) as the language model. BLANC Lita et al. (2005) measures how well a generated summary aids a pre-trained language model in performing a language understanding task on the original document. The score reflects the informativeness of the generated text. LENS Maddela et al. (2023) is a trainable metric which aims to reflect how closely the generated text aligns with human references in terms of content and style.
4.3.3 QA-based
QuestEval Scialom et al. (2021) combines scores obtained through the FEQA and SummaQA metrics. It employs a question-answering model to answer questions formed from the reference text (SummaQA) or the generated summary (FEQA), extracting answers from the opposite source. QuestEval evaluates factual accuracy, coherence, informativeness, and relevance.
4.4 Human Annotation
We adapted proven methodologies for a thorough annotation process Zhang et al. (2023). Four graduate students, aged 24 to 28, from diverse academic backgrounds (i.e., computer science, psychology, communication science), all well-versed in English and familiar with meeting summarization, were involved.
From the QMSum general-summary test set, we chose the 35 general-summary samples, each containing meeting transcripts, gold standard summaries, and model-generated summaries, resulting in a total of 175 distinct samples for annotation. Annotators identified challenges Section 3.1 in the transcript and errors Section 3.2 in the generated summaries. The evaluation used yes/no questions, such as: "Does the given summary omit crucial information or provide insufficient detail about salient points?" corresponding to the missing information error. They further rated the frequency of challenges and the impact of errors for the LED model, using Likert scales from rare occurrence or minimal impact (1) to frequent presence or high impact (5).
To ensure reliable and consistent annotation, we implemented several measures. We assessed inter-annotator agreement using Krippendorff’s alpha, achieving an average of 0.81 (see Tables 2 and 3). Before the annotation task, a pilot test was administered as training for our annotators and to refine our guidelines and point out new error types at the same time. Using the validation dataset, the annotators practiced on QMSum summaries produced by the SLED+BART model Ivgi et al. (2022). For the actual annotation task, we further used gold summaries that the annotators should be able to evaluate correctly, otherwise, their annotation would be rejected. Regular review meetings maintained consistency and alignment during the annotation process, and for ongoing quality control, annotators highlighted summary segments where they detected errors. Discrepancies were discussed, with an expert annotator on standby for complex issues.
5 Analysis
5.1 Linking Challenges and Error Types
Errors Encoder-Decoder Autoregressive
missing information discourse structure*, implicit context coreference*
redundancy speaker dynamics, contextual turn-taking*
implicit context*, decision dynamics*
spoken language*, speaker dynamics
contextual turn-taking, low info density
wrong references (none) spoken language**, contextual turn-taking*
low information density**
incorrect reasoning speaker dynamics, implicits context
low information density
hallucination contextual items, implicit context* coreference*, contextual turns*
implicit context
incoherence spoken language*, coreference*
contextual turn-takin*, decision dynamics*
coreference
linguistic inaccuracy coreference, low information density spoken language*, speaker dynamics**
decision dynamics*, low information density
structural disorganization spoken language**, speaker dynamics**
contextual turns*, decision dynamics
low information density
corefernce*, contextual turns
implicit context*, decision dynamics
Table 4: The linkage between challenges and errors. No asterisk indicates low correlation, * signifies mid correlation, and ** denotes high correlation.
We analyzed the relationship between challenges and frequently observed errors for encoder-decoder and autoregressive architectures, using Point-biserial correlation based on the human annotations from Section 4.4. The results are presented in Table 4. For encoder-decoder models, errors such as incoherence, structural disorganization, and redundancy arise from most challenges. These errors are strongly linked with challenges like spoken language and speaker dynamics, suggesting encoder-decoder models struggle with maintaining coherence in these complex challenges. Conversely, wrong references, linguistic inaccuracy, and hallucination errors show limited associations with the examined challenges. Specifically, the limited association of hallucination reinforces current understandings, suggesting that the underlying causes remain elusive Maynez et al. (2020). Notably, the implicit context challenge works as a proxy for hallucination. The minimal correlation of the wrong reference error, which rarely occurs due to the sentence structure in generated summaries, and the linguistic inaccuracy error indicate that challenges such as coreference are well-handled by these models.
Regarding specific challenges, discourse structure mildly correlates with the missing information error, suggesting that while models navigate the broader meeting structure well, they occasionally miss details within distinct phases. The low information density challenge weakly correlates with the incorrect reasoning error, revealing occasional difficulties in salient meeting details extraction. However, the contextual turn-taking challenge aligns more with errors like redundancy and incoherence, revealing issues in capturing the dynamics on a granular level within the different meeting phases.
For autoregressive models, patterns differ. Incorrect reasoning is now the rarest observed error. It shows the weakest correlation, while wrong references, redundancy, and linguistic inaccuracy errors are the most prevalent, aligning with challenges like low information and spoken language. This reflects the model’s design, as it presents a shift in redundancy and linguistic inaccuracies compared to encoder-decoder models, with redundancies manifesting as introductory sentences that are then repetitively echoed in subsequent sentences. Linguistic inaccuracies, on the other hand, often emerge as contextually ambiguous terms. Equally prominent but stronger correlated are now the missing information and structural disorganization errors, showing links to the coreference challenge and emphasizing models’ tendencies to list topics without proper context.
Notably, some negative correlations also emerge, suggesting specific challenges might decrease the likelihood of certain errors, such as for low information density challenge and missing information error. These novel observations in the field warrant further exploration.
5.2 Correlation of Automatic Metrics and Human Annotation
Though numerous and frequently applied, existing automatic metrics aren’t tailored to the intricacies of meeting summarization. We analyze their reactions to the diverse error types observable in meeting summaries by answering the following three research questions.
RQ1: How do automatic metrics correlate with human assessments?
MI Red WR IR Hal Inc LI SD
ROUGE-1 -0.40* 0.17 -0.07 -0.18 0.05 -0.30 -0.12 -0.41*
ROUGE-2 -0.20 0.10 0.07 -0.29 0.10 -0.22 -0.10 -0.35*
ROUGE-L -0.29 0.08 -0.04 -0.21 0.02 -0.21 -0.09 -0.46**
BLEU -0.20 0.08 -0.09 -0.08 0.32 -0.13 0.05 -0.26
METEOR -0.26 0.27 0.16 -0.23 0.08 -0.20 0.02 -0.38*
BERTScore (F) -0.16 0.22 0.09 -0.06 0.10 -0.02 -0.01 -0.26
PPL -0.10 -0.10 0.44** -0.32 -0.08 0.09 0.24 -0.17
BLANC (TS) -0.19 0.05 -0.13 -0.13 -0.13 -0.42* -0.09 -0.36*
LENS 0.01 -0.38* -0.17 0.20 0.01 0.03 -0.03 0.45**
QuestEval (F) 0.10 -0.05 0.16 0.02 0.26 0.00 -0.03 -0.28
Table 5: Point-biserial correlation between automatic metrics and annotated errors. * denotes significance at p $\leq$ 0.05 and ** at p $\leq$ 0.01. The top row shows abbreviated error types from Section 3.2 in order. A negative correlation indicates worsening metric scores with increasing error occurrence. The three highest absolute values are highlighted in bold. We present F-scores (F) for BERTScore and QuestEval, with BERTScores being further rescaled.
In Table 5, we present Point-biserial correlations between automatic metrics and expert-annotated errors, specifically for LED model-generated summaries due to their pronounced error variance. Our analysis highlights some overarching patterns: No metric consistently correlates highly with all error types, underscoring the complexity of meeting summarization and the absence of a universal evaluative metric that captures human judgment well Gao and Wan (2022). However, some metrics show trends aligning with human judgment through significant negative scores. Though not designed for structural coherence, several metrics show significant negative correlations with structural disorganization errors. This indicates that such temporal and logical disorganization breaks the n-gram sequences and semantic flow. This trend can be observed as the score of ROUGE-1 is vastly more significant than the ROUGE-2 and ROUGE-L scores, aligning with observations in dialogue summarization Gao and Wan (2022). In particular, the gaps, disjoint narratives, and incorrect statements from errors such as missing information, incoherence, and incorrect reasoning influence metric scores, as these events do not align with n-grames in the reference or shift meaning, thereby affecting the semantic alignment Lin (2004). Errors such as redundancy, wrong references, and hallucination remain less frequently detected in metric evaluation scores. This may stem from the fact that summaries with these errors, depending on how many words were altered, can still closely match the reference regarding n-gram overlap and semantic similarity.
As anticipated, count-based metrics directly compare generated and reference texts and are acutely responsive to missing information Lin (2004). ROUGE stands out, with ROUGE-1, in particular, correlating significantly with information omissions, aligning with prior dialogue summarization studies Gao and Wan (2022). Similarly, BLEU identifies summaries with critical point omissions. Despite not being designed for it, count-based metrics significantly correlate with the structural disorganization error, possibly arising from disrupted n-gram sequences due to reorganized content Banerjee and Lavie (2005). Moreover, ROUGE and Meteor are potential indicators for incorrect reasoning and incoherence.
Among model-based metrics, BERTScore exhibits sensitivity to missing information, an anticipated trend possibly stemming from the score reflecting this in the semantic and contextual differences between candidate and gold summary in its score Zhang et al. (2020b). However, its correlation with structural disorganization is less direct, hinting at the influence of sentence alignment and structural coherence. BLANC, designed for evaluating coherence and fluency, naturally correlates negatively with incoherence errors. The text simplification metric, LENS, effectively captures redundancy errors. Model-based metrics generally present milder correlations, possibly due to discrepancies between training data and meeting contexts Gao and Wan (2022). While they seem to not align as well with human judgment as count-based metrics, they do offer insights into specific errors not captured by count-based metrics.
The correlation table shows that no single metric predominantly captures all error types, with most correlations being weak to moderate. Examining Table 5, a combination of metrics like ROUGE, BLANC, and LENS could serve as a proxy for various errors due to their notable moderate correlations. Still, this approach warrants additional evaluation and score weighting.
MI Red WR IR Hal Inc LI SD
ROUGE-1 0.07 0.24 -0.02 -0.17 0.10 -0.30 -0.09 -0.42*
ROUGE-2 -0.05 0.15 0.06 -0.29 0.13 -0.23 -0.11 -0.36*
ROUGE-L 0.03 0.15 -0.05 -0.22 0.10 -0.21 -0.08 -0.44**
BLEU -0.11 0.08 -0.06 -0.09 0.35 -0.14 0.03 -0.31
METEOR -0.06 0.21 0.16 -0.27 0.13 -0.21 -0.05 -0.36*
BERTScore (F) -0.25 0.19 0.08 -0.10 0.15 -0.02 -0.04 -0.31
PPL -0.02 -0.13 0.42 -0.29 0.02 0.00 0.23 -0.13
BLANC -0.03 0.04 -0.07 -0.15 -0.11 -0.44 -0.13 -0.36*
LENS -0.14 -0.42* -0.26 0.22 -0.07 0.00 -0.06 0.40Ü
QuestEval (F) 0.13 0.04 0.18 0.05 0.34* -0.03 -0.07 -0.24
Table 6: Spearman correlation between metrics and annotated error impacts. * denotes significance at p $\leq$ 0.05 and ** at p $\leq$ 0.01. The top row lists abbreviated error types from Section 3.2 in sequence. A negative correlation implies declining metric scores with rising error instances. The three most pronounced absolute values are emphasized in bold. F-scores (F) are given for BERTScore and QuestEval, with BERTScores being further rescaled.
RQ2: Do automatic metrics mask errors?
We categorize "masking" in two ways: indifference to an error, seen as a near-zero correlation, or positive reinforcement of a mistake, seen as a positive correlation. Count-based metrics predominantly show near-zero or negative correlations with errors. This indicates that while they might not favour errors, they do not sufficiently penalize specific mistakes. Such behaviour is expected Saadany and Orasan (2021); Akter et al. (2022), especially for errors that these metrics weren’t explicitly designed to detect or for which they can only act as a proxy.
In contrast, while some metrics show their known struggle with specific error types through near-zero scores (e.g., BERTScore and linguistic inaccuracy Hanna and Bojar (2021)), several model-based and QA-based metrics occasionally manifest positive correlations. For instance, QuestEval rewards errors like missing information and hallucination, a counterintuitive outcome given its designed focus on factual accuracy. The finding indicates that hallucinated details subtly bypass metric detection, likely due to the generated questions lacking the rigour to spot hallucinations, which can be coupled to the gap between the pre-training and meeting data Gao and Wan (2022). Similarly, BERTScore, emphasizing semantic and syntactic similarities, may reward redundant yet correct details Hanna and Bojar (2021). In contrast, Perplexity favours incorrect references if they preserve linguistic coherence and fluency, thus aligning with the language model’s expectations and yielding a lower (better) score. Notably, LENS significantly correlates with the structural disorganization error. LENS, although designed for text simplification and primarily emphasizing word-level and intra-sentence simplification Maddela et al. (2023), appears to struggle with capturing broader logical and temporal structures, resulting in its significant correlation with the structural disorganization error.
RQ3: How do the metric scores vary with the severity of the error?
To further scrutinize the relationship between metric scores and error severity, we examined how these scores fluctuate by the impact of errors. The Spearman correlation trends are depicted in Table 6. Across different models and error severities, most correlations are negligible to weak, and many lack statistical significance, emphasizing the limitations of current metrics in discerning error severity in meeting summarization. BLEU and QuestEval display significant positive correlations with hallucination errors, underscoring their vulnerability to hallucinated content. While some metrics, such as BERTScore and LENS, showed sensitivity to missing information in RQ1, their precision in assessing the error severity remains limited. This might stem from the gap between the pre-training data and meeting transcript for the three metrics utilising language models, as the meeting transcript was provided as part of the input Gao and Wan (2022). LENS, being closely aligned with text simplification, penalizes redundancy, while Perplexity is particularly sensitive to wrong references and linguistic inaccuracies. These observations underscore that, while not explicitly designed for certain errors, some metrics are intrinsically responsive to them.
6 Related Work
Research on meeting summarization has typically addressed challenges and error types in isolation, leaving their interrelation largely unexplored. For the parent-domain dialogue summarization, challenges have been collected Chen and Yang (2020b). While we adapted these to our domain and corpus, we found them lacking the detail required for meeting contexts. Previous works, especially by Tang et al. (2022); Wang et al. (2022b), laid the groundwork for our error typologies. This foundation was enhanced with insights from diverse studies and feedback from our annotators. Further, while Chen and Yang (2020b) correlate challenges and errors in dialogue summarization, their findings didn’t entirely align with our needs, though their annotation approach informed ours. Although automatic metrics for meeting summaries are underexplored, similar studies exist for text summarization using the CNN/Dailymail dataset Fabbri et al. (2021) and for dialogue summarization via the SAMSum dataset Gao and Wan (2022). We built upon these works by selectively adopting their metrics and augmenting them with measures commonly reported in recent meetings and dialogue summarization studies.
7 Conclusion
We’ve crafted a resource suite dedicated to meeting summarization evaluation, encompassing domain-specific challenges, linked error types, their predictive correlations for a given model architecture with a known set of challenges, and expert annotations of a QMSum subset. Our analysis highlights the constraints and misalignments of current metrics with human judgment in discerning error nuances. Metrics working well for other summarization tasks either do not react to errors or cannot reflect the impact on the quality of their score. A composite metric might be more effective, although its formulation warrants further research. Recent advancements in the community have highlighted the potential of LLM-based metrics, particularly with chain-of-thought or tree-of-thought prompting techniques. Utilizing an LLM like GPT-4 or Llama 2, prompted with our detailed annotator guidelines and supplemented by a few examples, presents a promising approach. This method could be especially effective in detecting complex errors like structural disorganization and incorrect reasoning. In the future, we plan to augment the entire QMSum dataset with our annotations. Drawing inspiration from the Stanford Context Word Similarity dataset’s impact on word sense ambiguity Huang et al. (2012), we see an enriched QMSum offering deeper insights into potential challenges for new techniques. Further, investigating additional datasets (e.g., MeetingBank) and multi-lingual datasets (e.g., ELITR, FREDSum) is of interest to study our findings across languages. Considering the current evolution of techniques, we further plan to extend the work as a (dialogue) summarization benchmarking pipeline to better capture how upcoming techniques handle the challenges and how well new metrics capture the performance of techniques. We encourage the research community to join our initiative by contributing model outputs and introducing new metrics.
Limitations
Our study, while offering a comprehensive analysis of challenges and errors in meeting summarization, primarily focuses on English-language summaries. This linguistic focus may lead to variations in challenges and errors across languages. Some challenges, especially those subtle to human perception, might be overlooked, creating potential gaps in our annotations. Our reliance on Google Scholar, despite its broad coverage, has its criticisms, as it tends to favor citation counts Fagan (2017) and may include less reputable sources Beall (2016). The QMSum dataset, with its 35 samples, provides statistical significance but represents only a fraction of potential meeting types. Thus, findings may vary across different meeting contexts. The challenges we associate with QMSum are inferred from human annotations rather than directly tested, and our model selection, while reflective of the current landscape, doesn’t capture every variant. Our choice of encoder-decoder models is comprehensive, but we miss out on architectures like hierarchical models due to accessibility issues which hamper the comparability of the models. The array of autoregressive models is continually expanding, and our selection is based on the current standings on the Huggingface LLM Leaderboard. Our method of linking challenges to errors is holistic, yet it might dilute strong connections between them, as we did not test the linking with isolated challenges. Certain metrics, like LENS and QuestEval, may have biased scores since they use the meeting transcript as input and aren’t domain-trained. Lastly, our findings, rooted in the nuances of meeting summarization, might not seamlessly apply to broader summarization domains like dialogue summarization, given each domain’s distinctive traits.
Ethics Statement and Broader Impact
Our research abides by ethical guidelines for AI research and is committed to privacy, confidentiality, and intellectual property rights. We’ve ensured that the datasets in our study, publicly available, do not house sensitive or personal details. While our study leverages existing resources and generative models, it’s important to note that these models can possess biases and may occasionally generate summaries with distortions, biases, or inappropriate content Gooding (2022). To counteract this, we’ve configured our models to omit potentially harmful or unsafe content. While our research aims to enhance meeting summarization to benefit communication and productivity across sectors, we’re acutely aware of the ethical challenges posed by AI in this domain. Meeting summarization models must be wielded with respect to privacy and consent, especially when processing sensitive or confidential material. It’s paramount that these models neither violate privacy nor perpetuate harmful biases. As the field evolves, we stress the importance of maintaining these ethical considerations and encourage fellow researchers to uphold them, ensuring that AI advancements in meeting summarization are both beneficial and ethically grounded. An integral aspect of our ethical commitment is reflected in our approach to annotator recruitment and management. The team of annotators, consisting of interns, student assistants, and doctoral students, was meticulously selected through internal channels. This strategy was chosen to uphold a high standard of annotation quality—a quality we found challenging to guarantee through external platforms such as Amazon Mechanical Turk. Ensuring fair compensation, these annotators were remunerated in accordance with institutional guidelines for their respective positions. Further, flexibility in the annotation process was also a priority. Annotators had the freedom to choose their working times and environments to prevent fatigue from affecting their judgment.
8 Bibliographical References

### v2

@@
 Meeting summarization has become a critical task considering the increase in online interactions.
-While new techniques are introduced regularly, their evaluation uses metrics not designed to capture meeting-specific errors, undermining effective evaluation.
-This paper investigates what the frequently used automatic metrics capture and which errors they mask by correlating automatic metric scores with human evaluations across a broad error taxonomy.
-We commence with a comprehensive literature review on English meeting summarization to define key challenges like speaker dynamics and contextual turn-taking and error types such as missing information and linguistic inaccuracy, concepts previously loosely defined in the field.
-We examine the relationship between characteristic challenges and errors by using annotated transcripts and summaries from Transformer-based sequence-to-sequence and autoregressive models from the general summary QMSum dataset.
-Through experimental validation, we find that different model architectures respond variably to challenges in meeting transcripts, resulting in different pronounced links between challenges and errors.
-Current default-used metrics struggle to capture observable errors, showing weak to mid-correlations, while a third of the correlations show trends of error masking.
-Only a subset reacts accurately to specific errors, while most correlations show either unresponsiveness or failure to reflect the error’s impact on summary quality.
+Despite new techniques being proposed regularly, the evaluation of meeting summarization techniques relies on metrics not tailored to capture meeting-specific errors, leading to ineffective assessment.
+This paper explores what established automatic metrics capture and the errors they mask by correlating metric scores with human evaluations across a comprehensive error taxonomy.
+We start by reviewing the literature on English meeting summarization to identify key challenges, such as speaker dynamics and contextual turn-taking, and error types, including missing information and linguistic inaccuracy, concepts previously loosely defined in the field.
+We then examine the relationship between these challenges and errors using human-annotated transcripts and summaries from encoder-decoder-based and decoder-only Transformer models on the QMSum dataset.
+Experiments reveal that different model architectures respond variably to the challenges, resulting in distinct links between challenges and errors.
+Current established metrics struggle to capture the observable errors, showing weak to moderate correlations, with a third of the correlations indicating error masking.
+Only a subset of metrics accurately reacts to specific errors, while most correlations show either unresponsiveness or failure to reflect the error’s impact on summary quality.
 What’s under the hood: Investigating Automatic Metrics on Meeting Summarization
-Frederic Kirstein1,2, Jan Philip Wahle1, Terry Ruas1, Bela Gipp1 1Georg-August-Universität Göttingen, Germany 2kirstein@gipplab.org
+Frederic Kirstein1,2, Jan Philip Wahle1, Terry Ruas1, Bela Gipp1 1University of Göttingen, Germany 2kirstein@gipplab.org
 1 Introduction
-With the surge in remote collaboration, there’s an amplified need for proficient meeting summarization Mroz et al.
+The rise in remote collaboration has increased the need for effective meeting summarization Mroz et al.
 (2018); Pratama et al.
 (2020), beneficial for participants and non-attendees.
-Despite advancements in addressing current meeting challenges like contextual turn-taking and discourse structure, existing evaluation metrics often inadequately reflect these Rennard et al.
+However, established evaluation metrics do not fully capture the challenges of meeting transcripts such as contextual turn-taking and discourse structure Rennard et al.
 (2023); Kumar and Kabiri (2022).
-The widespread default use of the ROUGE metric Lin (2004) has shown limitations Liu and Liu (2008); Cohan and Goharian (2016); Fabbri et al.
+The commonly used ROUGE metric Lin (2004) has limitations in reflecting summary quality Liu and Liu (2008); Cohan and Goharian (2016); Fabbri et al.
+(2021), and newer metrics like BERTScore Zhang et al.
+(2020b) have not been thoroughly tested for meeting summarization Kirstein et al.
 (2024).
-While the research community acknowledges the potential of newer metrics, such as BERTScore Zhang et al.
-(2020b), their efficacy in meeting summarization remains uncharted.
-This is a crucial shortcoming, as the high abstraction level, low extraction rate, and the requirement for complex reasoning power of meeting summarization present new challenges to automatic evaluation metrics, which need to be explored Gao and Wan (2022).
-Our study aims to investigate how commonly used automatic metrics correlate with human annotations, exploring the inherent qualities these metrics encapsulate in their scores.
-To unify the fragmented understanding of challenges posed by meeting summarization and the typical errors that manifest when these challenges are unmet.
+Yet, this assessment would be crucial because meeting summarization comes with unique challenges, such as high abstraction, low extraction rate, and complex reasoning Gao and Wan (2022).
+This study examines how automatic metrics relate to human annotations for meeting summarization and what they actually measure in their scores.
+We aim to create a unified understanding of the challenges in meeting summarization and the errors that occur when these challenges are unmet.
 Using the QMSum dataset Zhong et al.
-(2021), we employ experts to annotate challenges in meeting transcripts and errors in summaries.
-Various models, including domain-standard encoder-decoder architectures and notable autoregressive models, generated these summaries.
-Our methodological approach enables us to draw correlations between the challenges and errors, the annotated errors, and a set of eight automatic metrics.
-The insights derived reveal shortcomings in the current automatic metrics employed for meeting summarization evaluation: Structural disorganization errors are often penalized, reflecting human assessments, whereas hallucination errors tend to be rewarded.
-While adept at penalizing omissions, the ROUGE Lin (2004) metric struggles to discern the nuances in error impact on quality.
-Strikingly, around a third of metrics-error combinations either overlook or reward errors.
-For instance, Perplexity leans towards incorrect references, and Lens tends to favour structural disorganization.
-This underscores a pressing demand for refined evaluation methods in meeting summarization.
+(2021), we have experts annotate the challenges in the transcripts and errors in automatically generated summaries using various models, including domain-standard encoder-decoder architectures and notable decoder-only models.
+This setup allows us to find connections between challenges, errors, and eight automatic metrics.
+The results show problems with current automatic metrics for evaluating meeting summarization.
+Structural disorganization errors are often penalized, matching human judgments, but hallucination errors are sometimes rewarded.
+ROUGE Lin (2004) struggles to distinguish the impact of different errors on quality, even though it is good at penalizing omissions.
+Surprisingly, about a third of the metrics-error combinations either ignore or reward errors.
+For example, Perplexity favors incorrect references, and Lens favors structural disorganization.
+These observations highlight the need for better evaluation methods in meeting summarization.
 Our contributions are threefold:
   • We conduct a comprehensive literature review to identify and unify the challenges specific to English meeting summarization and the types of errors that commonly occur in model-generated summaries.
-  • We build the first direct correlations between the intrinsic challenges and observable errors they induce for encoder-decoder and autoregressive model architectures, providing a framework for tracking their impact.
+  • We build the first direct correlations between the intrinsic challenges and observable errors they induce for encoder-decoder and decoder-only model architectures, providing a framework for tracking their impact.
   • We rigorously evaluate the efficacy of nine prevalent automatic metrics, uncovering what they accurately measure, neglect, and their sensitivity to varying error severities, all corroborated by human annotations.
-We will release the codebases, annotations and guidelines to aid continued research in this domain.
+The codebase, annotations, and guidelines are available on GitHub:
 2 Methodology
-We conducted a focused literature review to identify challenges and observable errors when failing these challenges in English meeting summarization, specifically targeting papers published after the seminal 2019 introduction of the transformer-based BART Lewis et al.
-(2020) model, which left a lasting mark in the field.
-This further allows us to target the challenges in meeting summarization with our survey as they emerge and pertain to the currently used Transformer-based models (e.g., coreference resolution and spoken language).
-Besides this temporal filter, we set inclusion criteria to prioritize relevance over historical comprehensiveness and consider works that either introduce novel methodologies or contribute to defining challenges, with a preference for those employing transformer architectures.
-Exclusion criteria eliminate tangential studies, highly specialized papers not addressing broader challenges, multi-lingual and multi-modal approaches, and duplicates.
-We considered publications from 2023 irrespective of their citation index, acknowledging their emergent influence.
-The survey thereby orients on the PRISMA checklist to prevent bias.
-Our primary database was Google Scholar , chosen for its comprehensive coverage of both extractive and abstractive methodologies and advanced query capabilities.
-While ACL Anthology , Web of Science , and DBLP were assessed, Google Scholar yielded a more nuanced spectrum of relevant literature, conforming to our review scope spanning from 2019 to September 2023.
-A two-stage search strategy was employed: The initial phase involved broad queries combining "meeting summarization" with adjunctive terms such as challenge, literature review, or survey, combined through a logical operation.
-Subsequent refinement employed challenge-related keywords and synonyms retrieved from the review literature, ensuring alignment with our research focus.
-Papers were subsequently ranked based on their Google Scholar metrics, with the top 100 from each query considered.
-The corpus thus aggregated served as the primary source for our analysis.
-From the initial pool of 300 papers, we rigorously applied our criteria to select 18 core articles addressing challenges directly and an additional 40 that discuss challenges and observable errors indirectly through methodological contributions.
-Each shortlisted paper underwent detailed data extraction to identify relevant challenges, methodologies, foundational models, and metrics, which were then collated in a structured format for cross-validation.
-This allowed us to construct a comprehensive view of current challenges and observable errors in English meeting summarization.
-3 Literature Review
+We conducted a focused literature review to identify challenges and observable errors when failing these in English abstractive meeting summarization.
+We target papers published between the introduction of the BART model Lewis et al.
+(2020) in 2019, which serves as the typical backbone model for the field, and 2024.
+This scope allows us to focus on challenges pertaining to currently used Transformer-based models.
+We set additional inclusion criteria to prioritize relevance.
+We consider works that introduce novel methodologies or contribute to defining challenges, with a preference for transformer architectures, and exclude only loosely related studies, multi-lingual, multi-modal approaches, non-abstractive, and non-Transformer-based approaches.
+We considered publications from 2024, irrespective of their citation index, acknowledging their emergent influence.
+The survey follows the PRISMA checklist Page et al.
+(2021) to prevent bias.
+Google Scholar is chosen as the primary database for its comprehensive coverage and advanced query capabilities, yielding more works conforming to our review scope than Web of Science and DBLP.
+Our search comprises two stages: The initial phase involves broader queries combining "meeting summarization" with adjunctive terms such as challenge, literature review, or survey, while the subsequent refinement uses challenge-related keywords and synonyms.
+Papers are ranked based on Google Scholar metrics, considering the top 100 from each query, and serve as primary source for our analysis.
+We apply our criteria to the initial pool of 300 papers to select 18 core articles directly addressing challenges and 40 additional papers discussing challenges and errors implicitly through methodological contributions.
+Each shortlisted paper undergoes detailed data extraction to identify relevant challenges, methodologies, foundational models, and metrics collated in a structured format for cross-validation.
+This approach allows us to construct a comprehensive view of current challenges and observable errors in English meeting summarization.
+3 Definitions
 3.1 Challenges
@@
 Handle colloquialisms, domain-specific terminology, and various forms of linguistic noise, such as false starts, repetitions, and filler words Koay et al.
-(2020); Kumar and Kabiri (2022); Feng et al.
-(2022); Antony et al.
-(2023); Zhang et al.
+(2020); Kumar and Kabiri (2022); Antony et al.
 (2023), which can appear due to the spoken language nature of meetings.
-This can affect the accuracy and clarity of the generated summaries.
+This challenge can affect the accuracy and clarity of the generated summaries.
 Speaker dynamics.
-Accurately distinguish and track different speakers, their utterances, and specific roles (e.g., project manager, applicant), particularly when roles are topic-dependent Gu et al.
+Accurately distinguish and track different speakers, their utterances, and specific roles (e.g., project manager, applicant), particularly when roles are topic-dependent Khalifa et al.
+(2021); Gu et al.
 (2022); Rennard et al.
-(2023); Shinde et al.
-(2022); Feng et al.
-(2022); Antony et al.
-(2023); Kumar and Kabiri (2022); Khalifa et al.
 (2023).
-Failing to do so can introduce biases and result in incomplete or flawed summaries leaving out noteworthy elements.
+Failing to do so can introduce biases and result in incomplete or flawed summaries, leaving out noteworthy elements.
 Coreference.
@@
 (2021).
-Inaccurate handling can lead to ambiguous or incomplete summaries lacking contextual richness.
+Inaccurate handling can lead to ambiguous or incomplete summaries lacking context.
 Discourse structure.
@@
 (2022), e.g., during an argumentation, or swiftly shift from one topic to another Shinde et al.
-(2022); Feng et al.
 (2022).
@@
 Capture the evolving local dynamics of a meeting as it progresses through multiple speaker turns.
-The task involves accurately processing shifts in discourse, complicated by interruptions, repetitions, and redundancies Ma et al.
-(2022); Gu et al.
-(2022); Zhang and Zhao (2021); Zhang et al.
-(2021); Shinde et al.
+The task involves accurately processing shifts in discourse complicated by interruptions, repetitions, and redundancies Ma et al.
+(2022); Zhang and Zhao (2021); Shinde et al.
 (2022).
@@
 (2021).
-This can affect the depth and level of salience in the resulting summary.
-Data scarcity. The lack of diverse, high-quality real-world scenario training samples hampers model training Rennard et al. (2023); Kumar and Kabiri (2022); Jacquenet et al. (2019).
+This challenge can affect the resulting summary’s depth and level of salience.
+Data scarcity. The lack of diverse, high-quality, real-world scenario training samples hampers model training Rennard et al. (2023); Kumar and Kabiri (2022); Jacquenet et al. (2019).
 Long transcripts.
-Process long transcripts, which can result from longer meetings, due to the quadratic computational cost of transformer-based models, leading to efficiency issues with increasing dialogue length Rennard et al.
-(2023); Kumar and Kabiri (2022); Shinde et al.
-(2022); Feng et al.
-(2022); Jacquenet et al.
+Process long transcripts, which can result from longer meetings due to the quadratic computational cost of transformer-based models, leading to efficiency issues with increasing dialogue length Kumar and Kabiri (2022); Feng et al.
 (2022); Zhang et al.
 (2021).
-Only processing a sub-part of a meeting might lead to the exclusion of salient information and therefore result in an incomplete or wrong summary.
+Only processing a sub-part of a meeting might exclude salient information and, therefore, result in an incomplete or incorrect summary.
 Heterogeneous meeting formats.
@@
 (2023).
-Failure to adapt can result in summaries that miss key points or include irrelevant information
+Failure to adapt can result in summaries that miss key points or include irrelevant information.
 3.2 Error Types
-Errors in summaries arise when challenges aren’t correctly addressed.
-We’ve collated error types from the literature into six principal categories, which were observed in the summaries generated during our experiments as detailed in Section 5.
-Missing information.
+Errors in summaries arise when challenges are not correctly addressed.
+We have collated error types from the literature into six principal categories, observed in the summaries generated during our experiments as detailed in Section 5.
+Missing information (MI).
 This error involves missing information from the meeting, such as significant decisions or actions Zou et al.
-(2023); Chen and Yang (2020a).
+(2023); Chen and Yang (2020).
 We consider total omission, where relevant topics or points discussed are entirely absent from the summary, and insufficient detail, where the summary mentions a salient topic but does not capture its depth or the detailed discussion from the original meeting.
-Redundancy.
-The summary contains repeated or redundant information, impacting brevity and clarity Chen and Yang (2020a).
-Such repetitions can manifest in different ways: through reiterated key points, overuse of individual words, or the duplication of entire phrases.
-Wrong references. The model misattributes statements, opinions, or actions to incorrect meeting participants or omits their mention altogether Chen and Yang (2020a).
-Incorrect reasoning. The model draws conclusions that are not supported by the discussions in the meeting Chen and Yang (2020a).
-Hallucination.
+Redundancy (Red).
+The summary contains repeated or redundant information, impacting brevity and clarity Chen and Yang (2020).
+Such repetitions can manifest in different ways: reiterated key points, overuse of individual words, or duplicating entire phrases.
+Wrong references (WR). The model misattributes statements, opinions, or actions to incorrect meeting participants or omits their mention altogether Chen and Yang (2020).
+Incorrect reasoning (IR). The model draws conclusions that are not supported by the discussions in the meeting Chen and Yang (2020).
+Hallucination (Hal).
 The model produces inconsistencies, such as incorrect dates, names, or locations, not aligned with the meeting content Wang et al.
-(2022a); Ji et al.
+(2022); Ji et al.
 (2023).
-This encompasses intrinsic hallucinations, introducing events not present in the original meeting or contradicting the input, and extrinsic hallucinations, misrepresenting actual events.
-Incoherence.
+This encompasses extrinsic hallucinations, introducing events not present in the original meeting or contradicting the input, and intrinsic hallucinations, misrepresenting actual events.
+Incoherence (Inc).
 The model generates summaries with disjointed logic or flow.
-This manifests as intra-sentence disconnections and inter-sentence inconsistencies, such as flawed progression and incorrect contrastive expressions Wang et al.
+This manifests as intra-sentence disconnections and inter-sentence inconsistencies, such as flawed progression and incorrect, contrastive expressions Wang et al.
 (2023).
-During our human annotation process, as detailed in Section 4.4, we identified two additional error categories.
-Notably, our review of the pertinent literature revealed that these categories have not been explicitly defined in existing studies in a way that matches our findings.
-Linguistic inaccuracy.
+We have identified two additional error categories during our human annotation process, as detailed in Section 4.3.
+Notably, our review of the pertinent literature reveals that these categories have not been explicitly defined in existing studies in a way that matches our findings.
+Linguistic inaccuracy (LI).
 The model uses inappropriate, incorrect, or ambiguous language or fails to capture unique linguistic styles.
-This error type spans issues from adopting unsuitable words from the source to grammar mistakes, as well as employing contextually unclear or ambiguous terms.
-Structural disorganization.
+This error type spans issues from adopting unsuitable words from the source to grammar mistakes and employing contextually unclear or ambiguous terms.
+Structural disorganization (SD).
 The model may produce summaries that misrepresent the original order or logic of the meeting’s discourse, misplacing topics or events.
 This error covers only when the order of stated events is wrong.
-The inclusion of non-existing events or exclusion of events does not count towards this error.
+Including non-existing events or excluding events does not count towards this error.
 4 Experimental Framework
 Dataset # Meetings # Turns # Speakers avg. Len. of Meet. avg. Len. of Sum.
-ICSI 137 535.6 4.0 6007.7 70.5
-AMI 59 819 6.3 13317.3 53.7
+AMI 137 535.6 4.0 6007.7 70.5
+ICSI 59 819 6.3 13317.3 53.7
 WPCP 36 207.7 34.1 13761.9 80.5
@@
 4.1 Models
-Our primary model is the Longformer Encoder Decoder (LED) Beltagy et al.
-(2020) given its recognized summarization task performance.
-LED employs a unique local+global sliding window attention in its encoder and full self-attention in its decoder, accommodating lengthy documents efficiently.
-For a comprehensive assessment of challenges and errors in the context of the common encoder-decoder architecture, we incorporated DialogLED Zhong et al.
+We use the Longformer Encoder Decoder (LED) Beltagy et al.
+(2020) as our primary model for its strong summarization performance.
+LED employs local+global sliding window attention in the encoder and full self-attention in the decoder, efficiently handling lengthy documents.
+To comprehensively assess challenges and errors in encoder-decoder architectures, we further consider DialogLED Zhong et al.
 (2022) and PEGASUS-X Phang et al.
 (2022).
-DialogLED extends LED with dialogue-centric pre-training.
-In contrast, PEGASUS-X enhances the original PEGASUS Zhang et al.
+DialogLED extends LED with dialogue-centric pre-training, while PEGASUS-X enhances PEGASUS Zhang et al.
 (2020a) for extended inputs using staggered block-local attention.
-These models have been finetuned on QMSum.
-We also ventured into large autoregressive language models for meeting summarization, leveraging GPT3.5 turbo via ChatGPT and Zephyr-7B-$\alpha$, a refined version of Mistral-7B-v0.1111https://huggingface.co/mistralai/Mistral-7B-v0.1.
-Mistral’s adaptations employ sliding window attention and have gained recent attention for outperforming Llama2 Touvron et al.
+All models are finetuned on the QMSum training subset (general summaries) for three epochs.
+We also explore large language models, using GPT3.5 turbo via ChatGPT and Zephyr-7B-$\alpha$, a refined version of Mistral-7B-v0.1  [^1] https://huggingface.co/mistralai/Mistral-7B-v0.1.
+Mistral’s adaptations employ sliding window attention, outperforming Llama2 Touvron et al.
 (2023).
-These models, exemplifying the trends in language models, excel in zero-shot summarization Bhaskar et al.
-(2023).
-To tackle context size limitations, we employed a chunk-based approach Bhaskar et al.
-(2023), using the prompt "Create a TL;DR of the following meeting chunk", inspired by a Microsoft guideline  [^2] https://www.microsoft.com/en-us/microsoft-365-life-hacks/organization/using-chatgpt-creating-meeting-agendas-minutes-notes.
+To handle context size limitations, we use a chunk-based approach Bhaskar et al.
+(2023) with the prompt "Create a TL;DR of the following meeting chunk," inspired by a Microsoft guideline  [^2] https://www.microsoft.com/en-us/microsoft-365-life-hacks/organization/using-chatgpt-creating-meeting-agendas-minutes-notes.
+LLMs are used with a zero-shot setup.
+For the experiments, the models are grouped into encoder-decoder models containing LED, DialogLED, and Pegasus-X, and the decoder-only models, i.e., GPT-3.5 turbo and Zephyr-7B-$\alpha$.
+This grouping is motivated by similar architecture and error distribution and frequency patterns as shown in Table 8.
 4.2 Datasets
-We employ QMSum Zhong et al.
-(2021), a specialized, widely established, and currently most comprehensive dataset for query-based multi-domain meeting summarization.
-It includes a diverse range of meeting transcripts from academic meetings (ICSI), product meetings (AMI), and committee meetings (Welsh Parliament, Canadian Parliament: WPCP).
-Detailed statistics are available in Table 1.
+We use as input samples from QMSum Zhong et al.
+(2021), an established dataset for query-based multi-domain meeting summarization.
+It includes transcripts from academic (ICSI), product (AMI), and committee meetings (Welsh and Canadian Parliament: WPCP).
 ICSI Janin et al.
-(2003) offers transcripts from informal research meetings, showcasing linguistic challenges such as missing words, sentence restarts, and fillers.
+(2003) offers informal research meetings with linguistic challenges.
 AMI Mccowan et al.
-(2005) provides structured meetings with predefined participant roles and agendas yet still captures natural dialogue dynamics like interruptions and non-verbal cues.
-WPCP presents formal, agenda-driven discussions from UK and Canada Parliament committee meetings, allowing for a more structured summarization task.
-SL SD Cor DS CTT IC DD LID overall
-challenges 0.84 0.77 0.68 0.79 0.73 0.90 0.87 0.78 0.79
+(2005) provides staged meetings with natural dialogue dynamics.
+WPCP presents formal, agenda-driven discussions from UK and Canada Parliament committee meetings.
+Detailed statistics are in Table 1.
+While there are other datasets out there in the context of meeting summarization, such as the frequently used MeetingBank Hu et al.
+(2023) and ELITR Nedoluzhko et al.
+(2022) datasets, we do not include these.
+MeetingBank is conceptually close to the WPCP meetings, therefore not enhancing the diversity of our data selection and potentially introducing an imbalance in meeting type.
+ELITR aims to produce meeting minutes with bullet points containing key insights and actions discussed.
+Therefore, it does not fit the abstractive summaries we aim to investigate.
+Spoken Language Speaker Dynamic Co- reference Discourse Structure Contextual Turn-Taking Implicit Context Data Scarcity Low Inf. Density Overall
+detection 0.84 0.77 0.68 0.79 0.73 0.90 0.87 0.78 0.79
 frequency 0.84 0.71 0.78 0.70 0.89 0.91 0.90 0.84 0.82
@@
 MI Red WR IR Hal Inc LI SD overall
-Encoder-Decorder 0.74 0.875 0.83 0.87 0.79 0.61 0.86 0.71 0.78
-impact (LED) 0.82 0.92 0.85 0.73 0.80 0.89 0.87 0.79 0.83
-Autoregressive 0.76 0.87 0.85 0.79 0.77 0.90 0.90 0.71 0.82
-Table 3: Krippendorff’s alpha for inter-annotator agreement on errors for encoder-decoder (i.e., LED, DialogLED) and autoregressive models (i.e., GPT-3.5, Zephyr-7B-$\alpha$).
+detection Encoder-Decorder 0.74 0.875 0.83 0.87 0.79 0.61 0.86 0.71 0.78
+Decoder-only 0.76 0.87 0.85 0.79 0.77 0.90 0.90 0.71 0.82
+error impact (LED) 0.82 0.92 0.85 0.73 0.80 0.89 0.87 0.79 0.83
+Table 3: Krippendorff’s alpha for inter-annotator agreement on errors for encoder-decoder (i.e., LED, DialogLED) and decoder-only models (i.e., GPT-3.5, Zephyr-7B-$\alpha$).
 For the LED model, the agreement on error impact is also reported.
@@
 4.3 Metrics
-For aligning human annotation with automatic metrics, we employ a suite of metrics spanning count-based, model-based, and QA-based methodologies, chosen based on their prevalent use in meeting and dialogue evaluation Gao and Wan (2022).
-4.3.1 Count-based
+We align human annotations with automatic metrics considering count-based, model-based, and QA-based methodologies, chosen based on their prevalent use in meeting and dialogue evaluation Gao and Wan (2022).
+Count-based.
 ROUGE Lin (2004) is the default-used metric suite that assesses the overlap between n-grams in generated summaries that appear in the reference.
@@
 METEOR Banerjee and Lavie (2005) builds on BLEU by accounting for synonyms, word stems, and sentence structure, offering a holistic assessment of lexical, syntactic, and semantic alignment in precision and recall.
-4.3.2 Model-based
+Model-based.
 BERTScore Zhang et al.
-(2020b) utilizes a pre-trained BERT model to measure the contextual similarity between the generated and reference texts, reflecting both semantic and syntactic similarity.
-We report the rescaled  [^3] https://github.com/Tiiiger/bert score/blob/master/journal/rescale baseline.md BERTScore-F score.
-Perplexity (PPL) measures the uncertainty of a language model in predicting each word in a given text, with lower scores denoting better predictions.
-It evaluates the model’s grasp of language structure and word occurrence patterns, thereby evaluating the quality and fluency of utterances.
+(2020b) measures the contextual similarity between generated and reference texts using a pre-trained BERT model, reflecting semantic and syntactic similarity  [^3] We report the rescaled BERTScore-F score:
+https://github.com/Tiiiger/bert score/blob/master/journal/rescale baseline.md
+.
+Perplexity (PPL) measures a language model’s uncertainty in predicting words, evaluating the quality and fluency of utterances.
 We use GPT-2 Radford et al.
@@
 BLANC Lita et al.
-(2005) measures how well a generated summary aids a pre-trained language model in performing a language understanding task on the original document.
-The score reflects the informativeness of the generated text.
+(2005) measures how well a generated summary aids a language model in understanding the original document, reflecting informativeness.
 LENS Maddela et al.
-(2023) is a trainable metric which aims to reflect how closely the generated text aligns with human references in terms of content and style.
-4.3.3 QA-based
+(2023) is a trainable metric that assesses the alignment of generated text with human references in content and style.
+While this metric is also explored for meeting summarization, it is noteworthy that it is based on RoBERTa, adopting its maximum context length of 512 tokens, making it technically less suitable for assessing meeting summaries considering transcripts.
+QA-based.
 QuestEval Scialom et al.
-(2021) combines scores obtained through the FEQA and SummaQA metrics.
-It employs a question-answering model to answer questions formed from the reference text (SummaQA) or the generated summary (FEQA), extracting answers from the opposite source.
-QuestEval evaluates factual accuracy, coherence, informativeness, and relevance.
-4.4 Human Annotation
-We adapted proven methodologies for a thorough annotation process Zhang et al.
-(2023).
-Four graduate students, aged 24 to 28, from diverse academic backgrounds (i.e., computer science, psychology, communication science), all well-versed in English and familiar with meeting summarization, were involved.
-From the QMSum general-summary test set, we chose the 35 general-summary samples, each containing meeting transcripts, gold standard summaries, and model-generated summaries, resulting in a total of 175 distinct samples for annotation.
-Annotators identified challenges Section 3.1 in the transcript and errors Section 3.2 in the generated summaries.
-The evaluation used yes/no questions, such as: "Does the given summary omit crucial information or provide insufficient detail about salient points?" corresponding to the missing information error.
-They further rated the frequency of challenges and the impact of errors for the LED model, using Likert scales from rare occurrence or minimal impact (1) to frequent presence or high impact (5).
-To ensure reliable and consistent annotation, we implemented several measures.
-We assessed inter-annotator agreement using Krippendorff’s alpha, achieving an average of 0.81 (see Tables 2 and 3).
-Before the annotation task, a pilot test was administered as training for our annotators and to refine our guidelines and point out new error types at the same time.
-Using the validation dataset, the annotators practiced on QMSum summaries produced by the SLED+BART model Ivgi et al.
-(2022).
-For the actual annotation task, we further used gold summaries that the annotators should be able to evaluate correctly, otherwise, their annotation would be rejected.
-Regular review meetings maintained consistency and alignment during the annotation process, and for ongoing quality control, annotators highlighted summary segments where they detected errors.
-Discrepancies were discussed, with an expert annotator on standby for complex issues.
+(2021) combines FEQA Durmus et al.
+(2020) and SummaQA Scialom et al.
+(2019) scores, using a question-answering model to answer questions formed from the reference text (SummaQA) or the generated summary (FEQA), extracting answers from the opposite source.
+QuestEval evaluates factuality, coherence, informativeness, and relevance.
+Human annotation.
+We adapt proven methodologies Zhang et al.
+(2023) for a thorough annotation process involving four graduate students from diverse backgrounds (i.e., computer science, psychology, communication science), all well-versed in English and familiar with meeting summarization.
+From the QMSum general-summary test set, we choose 35 general-summary samples, each containing meeting transcripts, gold summaries, and model-generated summaries, resulting in 175 distinct samples for annotation.
+Annotators identify challenges (Section 3.1) in the transcript and errors Section 3.2 in the generated summaries using yes/no questions such as: "Does the given summary omit crucial information or provide insufficient detail about salient points?" corresponding to the missing information error.
+The annotators further rate the frequency of challenges and the impact of errors for the LED model using Likert scales from rare occurrence or minimal impact (1) to frequent presence or high impact (5).
+To ensure reliability and consistency, we assess inter-annotator agreement using Krippendorff’s alpha, achieving an average of 0.81 (see Tables 2 and 3).
+A preliminary pilot test serves as annotator training and refinement of guidelines.
+Regular review meetings are held to maintain consistency.
+Annotators also highlight summary segments with errors.
+Discrepancies are discussed, and an expert annotator is available to discuss complex issues.
+A full presentation of the annotation process is stated in Appendix A. Details on annotated labels are shown in Appendix B.
 5 Analysis
 5.1 Linking challenges and error types
-Errors Encoder-Decoder Autoregressive
+Errors Encoder-Decoder Decoder-only
 missing information discourse structure*, implicit context coreference*
@@
 low information density**
-incorrect reasoning speaker dynamics, implicits context
+incorrect reasoning speaker dynamics, implicit context
 low information density
@@
 low information density
-corefernce*, contextual turns
+coreference*, contextual turns
 implicit context*, decision dynamics
-Table 4: The linkage between challenges and errors. No asterisk indicates low correlation, * signifies mid correlation, and ** denotes high correlation.
-We analyzed the relationship between challenges and frequently observed errors for encoder-decoder and autoregressive architectures, using Point-biserial correlation based on the human annotations from Section 4.4.
-The results are presented in Table 4.
-For encoder-decoder models, errors such as incoherence, structural disorganization, and redundancy arise from most challenges.
-These errors are strongly linked with challenges like spoken language and speaker dynamics, suggesting encoder-decoder models struggle with maintaining coherence in these complex challenges.
+Table 4: The linkage between challenges and errors. No asterisk indicates low correlation, * signifies mid correlation (p $\leq$ 0.05), and ** denotes high correlation (p $\leq$ 0.01).
+We analyze the relationship between challenges and frequently observed errors for encoder-decoder and decoder-only architectures using Point-biserial correlation based on human annotations from Section 4.3.
+Table 4 presents the results.
+Encoder-decoder models exhibit strong links between incoherence, structural disorganization, and redundancy errors and challenges like spoken language and speaker dynamics, suggesting struggles with maintaining coherence.
 Conversely, wrong references, linguistic inaccuracy, and hallucination errors show limited associations with the examined challenges.
@@
 The minimal correlation of the wrong reference error, which rarely occurs due to the sentence structure in generated summaries, and the linguistic inaccuracy error indicate that challenges such as coreference are well-handled by these models.
-Regarding specific challenges, discourse structure mildly correlates with the missing information error, suggesting that while models navigate the broader meeting structure well, they occasionally miss details within distinct phases.
-The low information density challenge weakly correlates with the incorrect reasoning error, revealing occasional difficulties in salient meeting details extraction.
-However, the contextual turn-taking challenge aligns more with errors like redundancy and incoherence, revealing issues in capturing the dynamics on a granular level within the different meeting phases.
-For autoregressive models, patterns differ.
-Incorrect reasoning is now the rarest observed error.
-It shows the weakest correlation, while wrong references, redundancy, and linguistic inaccuracy errors are the most prevalent, aligning with challenges like low information and spoken language.
-This reflects the model’s design, as it presents a shift in redundancy and linguistic inaccuracies compared to encoder-decoder models, with redundancies manifesting as introductory sentences that are then repetitively echoed in subsequent sentences.
-Linguistic inaccuracies, on the other hand, often emerge as contextually ambiguous terms.
-Equally prominent but stronger correlated are now the missing information and structural disorganization errors, showing links to the coreference challenge and emphasizing models’ tendencies to list topics without proper context.
-Notably, some negative correlations also emerge, suggesting specific challenges might decrease the likelihood of certain errors, such as for low information density challenge and missing information error.
-These novel observations in the field warrant further exploration.
+Regarding specific challenges, the discourse structure challenge correlates slightly with the missing information error, indicating occasionally missed details within distinct phases.
+The low information density challenge weakly correlates with the incorrect reasoning error, revealing difficulties in extracting salient details.
+The contextual turn-taking challenge aligns with redundancy and incoherence errors, suggesting issues in capturing dynamics on a granular level within the different meeting phases.
+Decoder-only models show different patterns, with incorrect reasoning being the rarest and weakest correlated error.
+Wrong references, redundancy, and linguistic inaccuracy errors are most prevalent, aligning with low information and spoken language challenges.
+For LLMs, redundancies manifest as repetitive introductory sentences, while linguistic inaccuracies emerge as contextually ambiguous terms.
+Missing information and structural disorganization errors are equally prominent and strongly correlated, linking to the coreference challenge and emphasizing models’ tendencies to list topics without proper context.
+Some negative correlations suggest specific challenges might decrease the likelihood of certain errors, such as for low information density challenge and missing information error, warranting further exploration.
 5.2 Correlation of automatic metrics and human annotation
-Though numerous and frequently applied, existing automatic metrics aren’t tailored to the intricacies of meeting summarization.
+Though numerous and frequently applied, existing automatic metrics are not tailored to the intricacies of meeting summarization.
 We analyze their reactions to the diverse error types observable in meeting summaries by answering the following three research questions.
@@
 QuestEval (F) 0.10 -0.05 0.16 0.02 0.26 0.00 -0.03 -0.28
-Table 5: Point-biserial correlation between automatic metrics and annotated errors.
+Table 5: Point-biserial correlation between automatic metrics and annotated errors on samples generated by LED.
 * denotes significance at p $\leq$ 0.05 and ** at p $\leq$ 0.01.
@@
 We present F-scores (F) for BERTScore and QuestEval, with BERTScores being further rescaled.
-In Table 5, we present Point-biserial correlations between automatic metrics and expert-annotated errors, specifically for LED model-generated summaries due to their pronounced error variance.
-Our analysis highlights some overarching patterns: No metric consistently correlates highly with all error types, underscoring the complexity of meeting summarization and the absence of a universal evaluative metric that captures human judgment well Gao and Wan (2022).
+Table 5 shows Point-biserial correlations between automatic metrics and expert-annotated errors for LED model-generated summaries.
+Our analysis reveals that no metric consistently correlates highly with all error types, underscoring the complexity of meeting summarization and the absence of a universal metric that captures human judgment well Gao and Wan (2022).
 However, some metrics show trends aligning with human judgment through significant negative scores.
-Though not designed for structural coherence, several metrics show significant negative correlations with structural disorganization errors.
-This indicates that such temporal and logical disorganization breaks the n-gram sequences and semantic flow.
-This trend can be observed as the score of ROUGE-1 is vastly more significant than the ROUGE-2 and ROUGE-L scores, aligning with observations in dialogue summarization Gao and Wan (2022).
-In particular, the gaps, disjoint narratives, and incorrect statements from errors such as missing information, incoherence, and incorrect reasoning influence metric scores, as these events do not align with n-grames in the reference or shift meaning, thereby affecting the semantic alignment Lin (2004).
-Errors such as redundancy, wrong references, and hallucination remain less frequently detected in metric evaluation scores.
-This may stem from the fact that summaries with these errors, depending on how many words were altered, can still closely match the reference regarding n-gram overlap and semantic similarity.
-As anticipated, count-based metrics directly compare generated and reference texts and are acutely responsive to missing information Lin (2004).
-ROUGE stands out, with ROUGE-1, in particular, correlating significantly with information omissions, aligning with prior dialogue summarization studies Gao and Wan (2022).
-Similarly, BLEU identifies summaries with critical point omissions.
-Despite not being designed for it, count-based metrics significantly correlate with the structural disorganization error, possibly arising from disrupted n-gram sequences due to reorganized content Banerjee and Lavie (2005).
-Moreover, ROUGE and Meteor are potential indicators for incorrect reasoning and incoherence.
-Among model-based metrics, BERTScore exhibits sensitivity to missing information, an anticipated trend possibly stemming from the score reflecting this in the semantic and contextual differences between candidate and gold summary in its score Zhang et al.
+Though not designed for structural coherence, several metrics show significant negative correlations with structural disorganization errors, indicating that temporal and logical disorganization breaks n-gram sequences and semantic flow.
+ROUGE-1 exhibits a more significant score than ROUGE-2 and ROUGE-L, aligning with observations in dialogue summarization Gao and Wan (2022).
+In particular, the gaps, disjoint narratives, and incorrect statements from errors such as missing information, incoherence, and incorrect reasoning influence metric scores by not aligning with reference n-grams or shifting meaning Lin (2004).
+Redundancy, wrong references, and hallucination errors remain less frequently detected, as summaries with these errors can still closely match the reference in terms of n-gram overlap and semantic similarity.
+As expected, count-based metrics are responsive to missing information Lin (2004), with ROUGE-1 correlating significantly with information omissions Gao and Wan (2022) and BLEU identifying summaries with key content omissions.
+Count-based metrics significantly correlate with structural disorganization errors though not designed to detect this error, possibly because of disrupted n-gram sequences due to reorganized content Banerjee and Lavie (2005).
+ROUGE and Meteor are potential indicators of incorrect reasoning and incoherence.
+Among model-based metrics, BERTScore exhibits sensitivity to missing information, reflecting semantic and contextual differences between candidate and gold summaries Zhang et al.
 (2020b).
 However, its correlation with structural disorganization is less direct, hinting at the influence of sentence alignment and structural coherence.
-BLANC, designed for evaluating coherence and fluency, naturally correlates negatively with incoherence errors.
-The text simplification metric, LENS, effectively captures redundancy errors.
-Model-based metrics generally present milder correlations, possibly due to discrepancies between training data and meeting contexts Gao and Wan (2022).
-While they seem to not align as well with human judgment as count-based metrics, they do offer insights into specific errors not captured by count-based metrics.
+BLANC, designed for evaluating coherence and fluency, correlates as expected negatively with incoherence.
+LENS effectively captures redundancy errors.
+Model-based metrics present milder correlations, possibly due to discrepancies between training data and meeting contexts Gao and Wan (2022).
+While these metrics seem to not align as well with human judgment as count-based metrics, they offer insights into errors not captured by count-based metrics.
 The correlation table shows that no single metric predominantly captures all error types, with most correlations being weak to moderate.
-Examining Table 5, a combination of metrics like ROUGE, BLANC, and LENS could serve as a proxy for various errors due to their notable moderate correlations.
-Still, this approach warrants additional evaluation and score weighting.
+Examining Table 5, a combination of metrics like ROUGE, BLANC, and LENS could serve as a proxy for various errors, but this approach warrants additional evaluation and score weighting.
 MI Red WR IR Hal Inc LI SD
@@
 ROUGE-L 0.03 0.15 -0.05 -0.22 0.10 -0.21 -0.08 -0.44**
-BLEU -0.11 0.08 -0.06 -0.09 0.35 -0.14 0.03 -0.31
+BLEU -0.11 0.08 -0.06 -0.09 0.35* -0.14 0.03 -0.31
 METEOR -0.06 0.21 0.16 -0.27 0.13 -0.21 -0.05 -0.36*
@@
 BLANC -0.03 0.04 -0.07 -0.15 -0.11 -0.44 -0.13 -0.36*
-LENS -0.14 -0.42* -0.26 0.22 -0.07 0.00 -0.06 0.40Ü
+LENS -0.14 -0.42* -0.26 0.22 -0.07 0.00 -0.06 0.40
 QuestEval (F) 0.13 0.04 0.18 0.05 0.34* -0.03 -0.07 -0.24
-Table 6: Spearman correlation between metrics and annotated error impacts.
+Table 6: Spearman correlation between metrics and annotated error impacts, using summaries generated by LED as base.
 * denotes significance at p $\leq$ 0.05 and ** at p $\leq$ 0.01.
@@
 RQ2: Do automatic metrics mask errors?
-We categorize "masking" in two ways: indifference to an error, seen as a near-zero correlation, or positive reinforcement of a mistake, seen as a positive correlation.
-Count-based metrics predominantly show near-zero or negative correlations with errors.
-This indicates that while they might not favour errors, they do not sufficiently penalize specific mistakes.
-Such behaviour is expected Saadany and Orasan (2021); Akter et al.
-(2022), especially for errors that these metrics weren’t explicitly designed to detect or for which they can only act as a proxy.
-In contrast, while some metrics show their known struggle with specific error types through near-zero scores (e.g., BERTScore and linguistic inaccuracy Hanna and Bojar (2021)), several model-based and QA-based metrics occasionally manifest positive correlations.
-For instance, QuestEval rewards errors like missing information and hallucination, a counterintuitive outcome given its designed focus on factual accuracy.
-The finding indicates that hallucinated details subtly bypass metric detection, likely due to the generated questions lacking the rigour to spot hallucinations, which can be coupled to the gap between the pre-training and meeting data Gao and Wan (2022).
-Similarly, BERTScore, emphasizing semantic and syntactic similarities, may reward redundant yet correct details Hanna and Bojar (2021).
-In contrast, Perplexity favours incorrect references if they preserve linguistic coherence and fluency, thus aligning with the language model’s expectations and yielding a lower (better) score.
-Notably, LENS significantly correlates with the structural disorganization error.
-LENS, although designed for text simplification and primarily emphasizing word-level and intra-sentence simplification Maddela et al.
-(2023), appears to struggle with capturing broader logical and temporal structures, resulting in its significant correlation with the structural disorganization error.
+We categorize "masking" as indifference to an error (near-zero correlation) or positive reinforcement of a mistake (positive correlation).
+Count-based metrics predominantly show near-zero or negative correlations with errors, indicating they might not sufficiently penalize specific mistakes.
+This behavior is expected Saadany and Orasan (2021); Akter et al.
+(2022), especially for errors these metrics were not explicitly designed to detect or for which they can only act as a proxy.
+From Table 5 using summaries generated by LED as a base, we observe that some metrics show their known struggle with specific error types through near-zero scores (e.g., BERTScore and linguistic inaccuracy Hanna and Bojar (2021)), model-based metrics occasionally manifest positive correlations.
+Perplexity favors incorrect references if they preserve linguistic coherence and fluency, aligning with the language model’s expectations and yielding a lower (better) score.
+LENS appears to struggle with capturing broader logical and temporal structures.
+Despite being designed for text simplification, it significantly correlates with the structural disorganization error and primarily emphasizes word-level and intra-sentence simplification Maddela et al.
+(2023).
+While QuestEval and BETScore do not exhibit significant correlations, they show noteworthy trends: QuestEval may reward errors like missing information and hallucination, a counterintuitive outcome given its focus on factual accuracy.
+The finding could indicate that hallucinated details likely bypass metric detection due to generated questions lacking the rigor to spot them, which can be linked to the gap between pre-training and meeting data Gao and Wan (2022).
+BERTScore, emphasizing semantic and syntactic similarities, may reward redundant yet correct details Hanna and Bojar (2021).
 RQ3: How do the metric scores vary with the severity of the error?
-To further scrutinize the relationship between metric scores and error severity, we examined how these scores fluctuate by the impact of errors.
-The Spearman correlation trends are depicted in Table 6.
-Across different models and error severities, most correlations are negligible to weak, and many lack statistical significance, emphasizing the limitations of current metrics in discerning error severity in meeting summarization.
+To identify the relationship between metric scores and error severity, we analyze how these scores fluctuate by the impact of errors.
+Table 6 shows the Spearman correlation trends, using summaries generated by LED as a base.
+Across different models and error severities, most correlations are negligible to weak, and many lack statistical significance.
+This observation emphasizes the limitations of current metrics in discerning error severity in meeting summarization.
 BLEU and QuestEval display significant positive correlations with hallucination errors, underscoring their vulnerability to hallucinated content.
-While some metrics, such as BERTScore and LENS, showed sensitivity to missing information in RQ1, their precision in assessing the error severity remains limited.
-This might stem from the gap between the pre-training data and meeting transcript for the three metrics utilising language models, as the meeting transcript was provided as part of the input Gao and Wan (2022).
-LENS, being closely aligned with text simplification, penalizes redundancy, while Perplexity is particularly sensitive to wrong references and linguistic inaccuracies.
-These observations underscore that, while not explicitly designed for certain errors, some metrics are intrinsically responsive to them.
+While metrics such as BERTScore and LENS show sensitivity to missing information in RQ1, their precision in assessing error severity is limited.
+This limitation may stem from the gap between training data and meeting transcripts for the metrics utilizing language models, as the meeting transcript was provided as part of the input Gao and Wan (2022).
+LENS, closely aligned with text simplification, penalizes redundancy.
+These observations underscore that some metrics are intrinsically responsive to specific errors, even if not explicitly designed for them.
+Perplexity does not exhibit significant correlations.
 6 Related Work
-Research on meeting summarization has typically addressed challenges and error types in isolation, leaving their interrelation largely unexplored.
-For the parent-domain dialogue summarization, challenges have been collected Chen and Yang (2020b).
-While we adapted these to our domain and corpus, we found them lacking the detail required for meeting contexts.
-Previous works, especially by Tang et al.
+Research on meeting summarization typically addresses challenges and error types in isolation, leaving their interrelation unexplored.
+For dialogue summarization, challenges are collected Chen and Yang (2020) but lack the detail required for meeting contexts.
+Previous works by Tang et al.
 (2022); Wang et al.
-(2022b), laid the groundwork for our error typologies.
-This foundation was enhanced with insights from diverse studies and feedback from our annotators.
-Further, while Chen and Yang (2020b) correlate challenges and errors in dialogue summarization, their findings didn’t entirely align with our needs, though their annotation approach informed ours.
-Although automatic metrics for meeting summaries are underexplored, similar studies exist for text summarization using the CNN/Dailymail dataset Fabbri et al.
-(2021) and for dialogue summarization via the SAMSum dataset Gao and Wan (2022).
-We built upon these works by selectively adopting their metrics and augmenting them with measures commonly reported in recent meetings and dialogue summarization studies.
+(2022) lay the groundwork for our error typologies, enhanced with insights from diverse studies and annotator feedback.
+Chen and Yang (2020) correlate challenges and errors in dialogue summarization, but their findings do not entirely transfer to English abstractive meeting summarization, though their annotation approach informs ours.
+Automatic metrics for meeting summaries are underexplored, but similar studies exist for text summarization using the CNN/Dailymail dataset Fabbri et al.
+(2021) and dialogue summarization via the SAMSum dataset Gao and Wan (2022).
+We build upon these works by selectively adopting their metrics and augmenting them with measures commonly reported in recent meetings and dialogue summarization studies.
 7 Conclusion
-We’ve crafted a resource suite dedicated to meeting summarization evaluation, encompassing domain-specific challenges, linked error types, their predictive correlations for a given model architecture with a known set of challenges, and expert annotations of a QMSum subset.
-Our analysis highlights the constraints and misalignments of current metrics with human judgment in discerning error nuances.
-Metrics working well for other summarization tasks either do not react to errors or cannot reflect the impact on the quality of their score.
-A composite metric might be more effective, although its formulation warrants further research.
-Recent advancements in the community have highlighted the potential of LLM-based metrics, particularly with chain-of-thought or tree-of-thought prompting techniques.
-Utilizing an LLM like GPT-4 or Llama 2, prompted with our detailed annotator guidelines and supplemented by a few examples, presents a promising approach.
-This method could be especially effective in detecting complex errors like structural disorganization and incorrect reasoning.
-In the future, we plan to augment the entire QMSum dataset with our annotations.
-Drawing inspiration from the Stanford Context Word Similarity dataset’s impact on word sense ambiguity Huang et al.
-(2012), we see an enriched QMSum offering deeper insights into potential challenges for new techniques.
-Further, investigating additional datasets (e.g., MeetingBank) and multi-lingual datasets (e.g., ELITR, FREDSum) is of interest to study our findings across languages.
-Considering the current evolution of techniques, we further plan to extend the work as a (dialogue) summarization benchmarking pipeline to better capture how upcoming techniques handle the challenges and how well new metrics capture the performance of techniques.
-We encourage the research community to join our initiative by contributing model outputs and introducing new metrics.
+We developed a resource suite for meeting summarization evaluation, covering domain-specific challenges, linked error types, predictive correlations for model architectures with known challenges, and expert annotations of the QMSum subset.
+Our analysis highlighted the limitations and misalignments of current metrics with human judgment in discerning error nuances.
+Metrics that work well for other summarization tasks either did not react to errors or cannot reflect the impact on quality in their scores.
+A composite metric may be more effective, but its formulation requires further research.
+Recent advancements highlighted the potential of LLM-based metrics.
+Utilizing an LLM prompted with our detailed annotator guidelines and supplemented by examples presents a promising approach for detecting complex errors like structural disorganization and incorrect reasoning.
+Given the current evolution of techniques, we plan to extend the work as a (dialogue) summarization benchmarking pipeline to capture better how upcoming techniques handle challenges and how well new metrics capture their performance.
+We encourage the research community to contribute model outputs and introduce new metrics to this initiative.
+Acknowledgements
+This work was supported by the Lower Saxony Ministry of Science and Culture and the VW Foundation. Frederic Kirstein was supported by the Mercedes-Benz AG Research and Development.
 Limitations
@@
 Our reliance on Google Scholar, despite its broad coverage, has its criticisms, as it tends to favor citation counts Fagan (2017) and may include less reputable sources Beall (2016).
-The QMSum dataset, with its 35 samples, provides statistical significance but represents only a fraction of potential meeting types.
-Thus, findings may vary across different meeting contexts.
-The challenges we associate with QMSum are inferred from human annotations rather than directly tested, and our model selection, while reflective of the current landscape, doesn’t capture every variant.
-Our choice of encoder-decoder models is comprehensive, but we miss out on architectures like hierarchical models due to accessibility issues which hamper the comparability of the models.
-The array of autoregressive models is continually expanding, and our selection is based on the current standings on the Huggingface LLM Leaderboard.
+The QMSum dataset, with its 35 samples, provides statistical significance but represents only a fraction of potential meeting types. Thus, findings may vary across different meeting contexts.
+We include a total of 175 samples, comparable to the original QMSum dataset (i.e., 232 samples), one of the most used datasets in meeting summarization.
+The considered dataset maintains a diverse representation of meeting types (academic, business, parliament) within the dataset.
+We acknowledge the limitation that QMSum might not cover all meeting types in existence.
+However, we focus on creating a unified understanding of the challenges in meeting summarization and the errors that occur when these challenges are unmet.
+For that, we used QMSum as a proxy.
+The challenges we associate with QMSum are inferred from human annotations rather than directly tested, and our model selection, while reflective of the current landscape, does not capture every variant.
+Our choice of encoder-decoder models is comprehensive, but we miss out on architectures like hierarchical models due to accessibility issues, which hamper the comparability of the models.
+The array of large language models is continually expanding, and our selection is based on the standings on the Huggingface LLM Leaderboard at the time of writing.
 Our method of linking challenges to errors is holistic, yet it might dilute strong connections between them, as we did not test the linking with isolated challenges.
-Certain metrics, like LENS and QuestEval, may have biased scores since they use the meeting transcript as input and aren’t domain-trained.
+Specific metrics, like LENS and QuestEval, may have biased scores since they use the meeting transcript as input and are not domain-trained.
 Lastly, our findings, rooted in the nuances of meeting summarization, might not seamlessly apply to broader summarization domains like dialogue summarization, given each domain’s distinctive traits.
@@
 Our research abides by ethical guidelines for AI research and is committed to privacy, confidentiality, and intellectual property rights.
-We’ve ensured that the datasets in our study, publicly available, do not house sensitive or personal details.
-While our study leverages existing resources and generative models, it’s important to note that these models can possess biases and may occasionally generate summaries with distortions, biases, or inappropriate content Gooding (2022).
-To counteract this, we’ve configured our models to omit potentially harmful or unsafe content.
-While our research aims to enhance meeting summarization to benefit communication and productivity across sectors, we’re acutely aware of the ethical challenges posed by AI in this domain.
+We have ensured that the datasets in our study, which are publicly available, do not house sensitive or personal details.
+While our study leverages existing resources and generative models, it is important to note that these models can possess biases and may occasionally generate summaries with distortions, biases, or inappropriate content Gooding (2022).
+We have configured our models to omit potentially harmful or unsafe content to counteract this.
+While our research aims to enhance meeting summarization to benefit communication and productivity across sectors, we are acutely aware of the ethical challenges posed by AI in this domain.
 Meeting summarization models must be wielded with respect to privacy and consent, especially when processing sensitive or confidential material.
@@
 This strategy was chosen to uphold a high standard of annotation quality—a quality we found challenging to guarantee through external platforms such as Amazon Mechanical Turk.
-Ensuring fair compensation, these annotators were remunerated in accordance with institutional guidelines for their respective positions.
+Ensuring fair compensation, these annotators were reimbursed following institutional guidelines for their respective positions.
 Further, flexibility in the annotation process was also a priority.
-Annotators had the freedom to choose their working times and environments to prevent fatigue from affecting their judgment.
-8 Bibliographical References
+Annotators were free to choose their working times and environments to prevent fatigue from affecting their judgment.
+Appendix A Annotation process details
+Following, we describe the details of our annotation process.
+Annotator selection:
+Our annotation team comprises four graduate students, officially employed as interns or doctoral candidates through standardized contracts.
+We select them from a pool of volunteers based on their availability to complete the task without time pressure and their English proficiency (native speakers or C1-C2 certified).
+This ensures they can comprehend meeting transcripts, human-written gold summaries from QMSum, and model-generated summaries.
+We aimed for gender balance (two male, two female) and diverse backgrounds, resulting in a team of two computer science students, one psychology student, and one communication science student, aged 24-28.
+Preparation:
+We prepare a comprehensive handbook for our annotators, detailing the project context and defining challenges and error types.
+Each definition includes two examples: one with minimal impact (e.g., slight information redundancy) and one with high impact (e.g., repeated information throughout).
+The handbook explains the two-part rating system: a binary yes/no for the existence of a challenge or error, followed by a 1-5 Likert scale impact/frequency rating if a characteristic is observed as existing.
+This impact/frequency scoring is only used for the challenges (frequency, 1 low - 5 high) and errors (impact, 1 low - 5 high) produced by the LED model.
+Annotators are further tasked to provide reasoning for each decision.
+The handbook does not specify an order for processing errors or challenges.
+We provide the handbook in English and the annotators’ native languages, using professional translations.
+The handbook could be used throughout the whole annotation process as a reference.
+We set up a five-week timeline for the annotation process, preceded by a one-week onboarding period.
+The first two weeks feature twice-weekly check-ins with annotators, which are reduced to weekly meetings for the following three weeks.
+Separate quality checks without the annotators are scheduled weekly.
+(Note: week refers to a regular working week)
+Onboarding:
+The onboarding week is dedicated to getting to know the project and familiarizing with the definitions and data.
+We begin with a kick-off meeting to introduce the project and explain the handbook, mainly focusing on each definition.
+We note initial questions to revise the handbook potentially.
+Annotators are provided with 35 samples generated by SLED+BART Ivgi et al.
+(2022), chosen for their balance of identifiable errors and good-quality summaries while capable of processing the whole meeting.
+After the first 15 samples, we hold individual meetings to clarify any confusion and update the guidelines accordingly.
+The remaining 20 samples are then annotated using these updated guidelines.
+A second group meeting this week addresses any new definitional issues.
+After the group meeting, we meet individually with annotators to review their work, ensuring quality and understanding of the task and samples.
+All four annotators demonstrate reliable performance and good comprehension of the task and definitions, judging from the reasoning they provided for each decision and annotation.
+We computed an inter-annotator agreement score using Krippendorff’s alpha, achieving 0.81, indicating sufficiently high overlap.
+Annotation Process:
+We continue the annotation process similarly.
+Each week, we distribute all 35 samples generated by one model to one of the annotators.
+Consequently, one annotator works through all samples of one model in one week.
+After five weeks, all samples have been processed by all annotators.
+Annotators are unaware of the summary-generating model and are given a week to complete their set at their own pace and break times.
+Quiet working rooms were provided if needed for concentration.
+To mitigate position bias, the sample order is randomized for each annotator.
+Annotators can choose their annotation order for each sample and are allowed to revisit previous samples to adapt ratings.
+To simplify the process, we frame each error type as a question, such as "Does the summary omit crucial information or provide insufficient detail about salient points?" for the missing information error.
+Regular meetings are held to address questions on definitions or emerging issues.
+During the quality checks the authors perform, we look for incomplete annotations, missing explanations, and signs of misunderstanding judging from the provided reasoning.
+If we find such a lack of quality, the respective annotator will be notified to re-do the annotation.
+The quality checks are not used to bias annotators in their ratings but to ensure a complete and consistent dataset.
+After the five weeks, we compute inter-annotator agreement scores (shown in Tables 2 and 3).
+In case we observe a significant difference across annotators at this point, we have planned a dedicated meeting to discuss such cases with all annotators and a senior annotator to ensure that an understanding issue of the task or definition did not lead to the different ratings.
+Annotators spend 43 minutes per sample, completing about seven samples daily.
+Handling of scheduling conflicts:
+Given that our annotators have other commitments, we anticipate potential scheduling conflicts.
+We allow flexibility for annotators to complete their samples beyond the week limit if needed, reserving a sixth week as a buffer.
+Despite these provisions, all annotators completed their assigned samples within the original weekly timeframes.
+Appendix B Statistics on Annotation Labels
+In Tables 7 and 8, we report statistics of the annotated dataset, showing how many are showing the respective challenge and error type.
+Spoken Language Speaker Dynamic Co- reference Discourse Structure Contextual Turn-Taking Implicit Context Data Scarcity Low Inf. Density
+# detections 32 (91%) 35 (100%) 35 (100%) 34 (97%) 33 (94%) 5 (14%) 30 (86%) 35 (100%)
+Table 7: Statistics on the occurance of different challenges in the QMSum samples used as input. We report total number and corresponding percentage.
+MI Red WR IR Hal Inc LI SD LED 31 (89%) 14 (40%) 4 (11%) 7 (20%) 9 (26%) 13 (37%) 3 (9%) 20 (57%) DialogLED 33 (94%) 18 (51%) 4 (11%) 8 (23%) 8 (23%) 18 (51%) 3 (9%) 22 (63%) Pegasus-X 34 (97%) 27 (77%) 14 (40%) 16 (46%) 13 (37%) 23 (66%) 10 (29%) 28 (80%) GPT3.5 34 (97%) 7 (20%) 3 (9%) 1 (3%) 5 (14%) 9 (26%) 9 (26%) 22 (63%) Zephyr 34 (97%) 7 (20%) 3 (9%) 1 (3%) 8 (23%) 11 (31%) 11 (31%) 22 (63%)
+Table 8: Statistics on the occurance of different error types in the model-generated summaries. We report total number and corresponding percentage.
+Citation for this Paper